# Adaptive Rank Allocation for Federated Parameter- Efficient Fine-Tuning of Language Models

### v2

@@ Abstract
Pre-trained Language Models (PLMs) have demonstrated their superiority and versatility in modern Natural Language Processing (NLP), effectively adapting to various downstream tasks through further fine-tuning. Federated Parameter-Efficient Fine-Tuning (FedPEFT) has emerged as a promising solution to address privacy and efficiency challenges in distributed training for PLMs on resource-constrained local devices. However, our measurements reveal two key limitations of FedPEFT: heterogeneous data across devices leads to significant performance degradation, and a fixed parameter configuration results in communication inefficiency. To overcome these limitations, we propose FedARA, a novel Adaptive Rank Allocation framework for federated parameter-efficient fine-tuning of language models. Specifically, FedARA employs truncated Singular Value Decomposition (SVD) adaptation to enhance similar feature representation across clients, significantly mitigating the adverse effects of data heterogeneity. Subsequently, it utilizes dynamic rank allocation to progressively identify critical ranks, effectively improving communication efficiency. Lastly, it leverages rank-based module pruning to automatically remove inactive modules, steadily reducing local computational cost and memory usage in each federated learning round. Extensive experiments show that FedARA consistently outperforms baselines by an average of 6.95% to 8.49% across various datasets and models under heterogeneous data while significantly improving communication efficiency by 2.40$\times$. Moreover, experiments on various edge devices demonstrate substantial decreases in total training time and energy consumption by up to 48.90% and 46.95%, respectively.
@@ I Introduction
Pre-trained Language Models (PLMs) [1, 2, 3] have demonstrated extraordinary performance and foundational capabilities in modern Natural Language Processing (NLP). Fine-tuning of PLMs further enhances their performance on downstream tasks, such as sequence classification, question answering, and causal language modeling. Pre-training and fine-tuning are evolving into two distinct yet mutually reinforcing stages in NLP. 1) Pre-training: it involves self-supervised training on large-scale corpora from sources like books, news articles, and Wikipedia [4], aiming to learn general representations of contextual language and create a generalizable pre-trained model. 2) Fine-tuning: it extends the pre-trained model to specific downstream tasks through supervised training, optimizing its performance to meet domain requirements. Overall, these two stages greatly enhance the model’s generality and flexibility across diverse NLP tasks.
@@
However, the massive scale of PLMs, often with hundreds of millions to billions of parameters [10], incurs substantial computation and communication costs in FL, rendering deployment on resource-constrained devices nearly impossible. To address this issue, a promising solution is to adopt Parameter-Efficient Fine-Tuning (PEFT) methods, which train only a minimal set of parameters using adaptable modules while keeping the pre-trained model frozen [11, 12, 13, 14, 15]. Among these PEFT methods, Low-Rank Adaptation (LoRA) [15] has gained widespread adoption for its effectiveness and efficiency, achieved by introducing two trainable low-rank matrices alongside the pre-trained model for diverse task adaptation. By combining PEFT with FL, distributed data resources on local devices can be fully leveraged for collaborative fine-tuning in a privacy-preserving manner. Consequently, Federated Parameter-Efficient Fine-Tuning (FedPEFT) shows significant advancements and finds wide-ranging applications across various domains [16, 17, 18, 19, 20, 21, 22, 23, 24]. Building on this progress, FedPEFT has increasingly gained recognition as a key method for distributed fine-tuning on local devices [25, 26]. Figure 1 shows the complete FedPEFT training process with LoRA adaptation in local clients.
Nevertheless, FedPEFT still faces two major challenges when applied to local devices. 1) FedPEFT exacerbates the impact of non-independent and identically distributed (non-IID) data on model performance. Due to the diverse preferences of local device users, data on these devices is typically non-IID. While performance degradation caused by non-IID data is a well-known challenge in FL, our preliminary measurements show that FedPEFT exacerbates this issue significantly. Specifically, on the 20News [27] dataset with the DistilBERT [3] model, different FedPEFT methods under non-IID data result in an average accuracy drop of 14.62% compared to IID settings. This significantly exceeds the roughly 6% drop observed with federated full fine-tuning [20]. Recently, several studies have also recognized this issue [20, 21], but their efforts have primarily focused on improving data initialization to provide a better starting point for training. Consequently, an effective solution to mitigate the impact of non-IID data on FedPEFT remains lacking.
2) Fixed parameter configurations in FedPEFT cause inefficient communication during FL. In FedPEFT, low-rank parameters can be placed at different positions within the model, including different layers or components like query, key, and value matrices. Current research often uses a fixed parameter configuration that assigns a uniform parameter size to all inserted modules, with the total parameter size serving as the overall parameter budget. However, our preliminary experiments reveal that the contribution of parameters at different positions to performance varies across models and datasets, with accuracy differences between configurations reaching up to 9.76%, indicating that there is no one-size-fits-all configuration for PEFT modules. Fixed configurations allocate the same parameter budget to modules with varying contributions to performance. The less impactful parameters are repeatedly transmitted over FL rounds without proportional performance benefits, which leads to communication inefficiency. Therefore, dynamic parameter allocation becomes especially crucial when the overall parameter budget is limited, particularly for resource-constrained devices.
@@
To address these challenges, we propose Adaptive Rank Allocation for Federated PEFT of language models (FedARA), designed to enhance performance and communication efficiency with minimized resource costs by integrating three novel methods: 1) Truncated SVD adaptation (Section IV-A). We hypothesize that the performance degradation of LoRA arises from its inherent structure and asymmetric initialization. Inspired by truncated Singular Value Decomposition (SVD), we propose a novel truncated SVD adaptation scheme that introduces an additional diagonal matrix into the low-rank subspace. This matrix enables independent feature scaling while preserving a symmetric initialization. It improves the ability to represent similar features across different clients, significantly alleviating the negative effects of non-IID data. 2) Dynamic rank allocation (Section IV-B). To avoid fixed parameter configurations, we present a dynamic rank allocation method that generates local rank masks through triplets on each client and produces global rank masks via threshold-based arbitration on the server. During federated fine-tuning, less critical ranks are progressively pruned, retaining only significant ones based on the global rank masks, effectively improving the communication efficiency of FedPEFT. 3) Rank-based module pruning (Section IV-C). Building on the above two methods, the rank of SVD modules dynamically evolves over FL rounds, we thereby adopt a rank-based module pruning mechanism to automatically remove inactive modules during training. By continuously monitoring rank changes, modules with ranks that fall to zero are safely excluded from the trainable parameters, steadily reducing the computational and storage overhead for client-side training.
@@
Results. We implemented FedARA and evaluated it on three different types of advanced edge devices, NVIDIA Jetson AGX Orin, Orin Nano, and Raspberry Pi 5, which have hardware specifications similar to common resource-constrained devices. We utilize two pre-trained models across four diverse NLP datasets. FedARA outperforms state-of-the-art (SOTA) methods [24, 28, 29] by an average of 6.95% to 8.49% under non-IID data while significantly improving communication efficiency by an average of 2.40$\times$. Remarkably, the total training time on the above devices also decrease by 27.65%, 46.57%, and 48.90%, respectively, and the energy consumption drops by 46.95% on Orin Nano. Ablation experiments reveal that truncated SVD adaptation improves the average accuracy by 7.71%, dynamic rank allocation reduces communication overhead by 70.82% per round while maintaining on-par performance, and rank-based module pruning lowers the average local computation and GPU memory usage per round by 10.81% and 31.67%, respectively.
@@
  • We create and integrate three innovative methods into FedARA: 1) truncated SVD adaptation strengthens the structure and adjusts initialization to enhance similar feature representation across clients; 2) dynamic rank allocation generates rank masks locally and arbitrates them globally to reduce the communication of less important parameters; 3) rank-based module pruning monitors dynamic rank changes to automatically remove inactive modules.
@@ II-A Federated Parameter-Efficient Fine-Tuning (FedPEFT)
FedPEFT orchestrates PEFT strategies within the FL framework to efficiently update local models, drastically reducing communication overhead by only transmitting a small number of PEFT modules. These PEFT modules can be replaced by Adapter [11, 12], Prompt[13, 14], or LoRA [15]. Among them, LoRA-based methods are especially popular because they introduce no additional latency during inference. FedPEFT not only preserves privacy but also effectively aggregates diverse knowledge across clients, facilitating the collaborative development of a high-quality global model. By harnessing the advantages of both PEFT and FL, FedPEFT has become an essential technique for deploying fine-tuning on resource-constrained devices [25, 26, 21, 19, 17, 22, 23].
@@ II-B Preliminary Measurements
Observation-1: FedPEFT is more effective than FedFFT on advanced edge devices. FedPEFT comprises two main methods: adapter-based [25] and LoRA-based methods [16, 17, 19, 20, 22, 23]. LoRA-based methods have gradually replaced adapter-based methods to avoid additional inference latency. According to established work [25], FedPEFT achieves comparable accuracy to federated full fine-tuning (FedFFT) while significantly reducing local training time by 1.6$\times$, communication cost by 200$\times$, and memory usage by 1.5$\times$ on previous-generation edge devices like NVIDIA Jetson TX2, Nano, and Raspberry Pi 4B. However, its impact on overall training time on advanced devices remains underexplored. To fill this gap, we evaluate FedLoRA on the 20News [27] dataset with the BERT [2] model using latest edge devices. Following FedPEFT’s typical settings [25], we set the communication bandwidth to 1MB/s. As depicted in Figure 2a, the results indicate that FedFFT’s overall training time is 42$\times$, 40$\times$, and 18$\times$ longer than FedLoRA on AGX Orin, Orin Nano, and RPi 5, respectively. FedPEFT demonstrates its effectiveness compared to FedFFT for cutting-edge embedded devices.
@@
Observation-3: The contribution of PEFT modules in FedPEFT to performance varies their inserted positions, leading to inefficient communication in fixed parameter configurations. In FedPEFT, low-rank modules can be placed at various layers and components, such as Query (Q), Key (K), Value (V), Out (O), F1, and F2. Here, O denotes the linear layer in the multi-head attention block, while F1 and F2 represent the two linear layers of the feed-forward network. As shown in Figure 2c, we allocated a fixed rank budget to specific components or layers to evaluate their impact on accuracy. The left sub-figure shows the allocation of the entire rank budget to a single layer, while the right sub-figure illustrates its distribution across a single component in all layers. The findings reveal that deeper layers near the output often have a more significant effect than shallower ones, with components F1 and F2 typically exerting a greater influence than component O. Moreover, the impact of these inserted positions varies across datasets, suggesting that there is no one-size-fits-all configuration. Fixed rank configurations can result in inefficient communication, as they repeatedly transmit less impactful parameters over hundreds of rounds, yielding minimal performance gains under a limited parameter budget.
@@
Insights: These observations underscore key directions for advancing our research. First, given its established advantages, such as improved training efficiency, future research should shift from comparative studies with FedFFT to addressing its unique challenges in FL. Second, the performance degradation of different PEFT methods under non-IID remains a critical barrier, necessitating further investigation into effective approaches. Third, fixed parameter configurations lead to communication inefficiency by repeatedly transmitting less impactful parameters over multiple FL rounds, particularly in resource-constrained environments. Finally, optimizing FedPEFT for edge devices requires further refinement, as different hardware presents distinct bottlenecks. To bridge these gaps, enhancing fine-tuning structures and dynamically adjusting parameters could be essential while simultaneously minimizing both communication and computational overhead.
@@ III Overview of FedARA
11:     $\{\hat{P}_{l+1}\},\{M_{l+1}\}\leftarrow\textsc{CLIENT\_TRAIN}(r,\hat{P}_{g},M_{g})$
@@ IV Design Details of FedARA
In this section, we provide a detailed description of the design of FedARA with the three closely correlated methods, including truncated SVD adaptation, dynamic rank allocation, and rank-based module pruning:
@@ IV-A Truncated SVD Adaptation
$W=W_{pre}+\Delta W=W_{pre}+BA,$ (1)
where $W_{pre},\Delta W\in\mathbb{R}^{d_{out}\times d_{in}}$, $B\in\mathbb{R}^{d_{out}\times r}$, and $A\in\mathbb{R}^{r\times d_{in}}$ with $r\ll\{d_{out},d_{in}\}$. Pre-trained weights $W_{pre}$ are frozen, while only the low-rank matrices B and A undergo fine-tuning. LoRA improves parameter efficiency compared to full fine-tuning when $r(d_{out}+d_{in})<d_{out}d_{in}$. The matrix $A$ is initialized with random Gaussian values, and $B$ is set to zero, ensuring $\Delta W=0$ at the beginning.
@@
To address this issue, we draw inspiration from DoRA [31], which decomposes pre-trained weights into magnitude and direction to imitate the behavior of full fine-tuning. However, its directional normalization and additional magnitude parameters introduce substantial computational and memory overhead, making it unsuitable for resource-constrained devices. Instead, inspired by truncated Singular Value Decomposition (SVD), which decomposes a matrix into the eigenvalues and eigenvectors of its low-rank subspace, we propose truncated SVD adaptation, formulated as
$W=W_{\text{pre}}+\Delta W\;=\;W_{\text{pre}}\;+\;BEA,$ (2)
where $E\in\mathbb{R}^{r\times r}$ is a diagonal matrix, and $B\in\mathbb{R}^{d_{\text{out}}\times r}$, $A\in\mathbb{R}^{r\times d_{\text{in}}}$. Matrices $A$ and $B$ are initialized with random Gaussian values, while $E$ is set to zero, ensuring $\Delta W=0$ at the start of training. Unlike the original truncated SVD, we do not enforce strict orthogonality constraints on $B$ and $A$, as such constraints limit adaptability and introduce additional computational overhead per round.
@@
As illustrated in Figure 4, a diagonal matrix $E$ is introduced between the low-rank subspace at a negligible cost ($r\ll\min(d_{\text{in}},d_{\text{out}})$), enabling the additional structural constraint while preserving a symmetric initialization. Specifically, the diagonal structure of the $E$ matrix imposes additional constraints, restricting updates to independent scaling of each latent dimension, which mitigates inter-client differences under non-IID data. Meanwhile, symmetric initialization prevents imbalanced gradient flow, providing clients with a more stable training starting point and promoting global consistency in federated fine-tuning under data heterogeneity. Interestingly, setting $E$ as the identity matrix reduces truncated SVD to standard LoRA, unifying truncated SVD and LoRA adaptation under the same umbrella.
Next, we analyze LoRA and truncated SVD adaptation from the perspective of the generalization bound [28]. Let $\mathbf{W}=\{W_{i}\}_{i=1}^{L}$ denote the $L$ parameter matrices of a pre-trained model. Define a subset $\mathcal{I}\subseteq\{1,2,\ldots,L\}$ as the indices of parameter matrices to be fine-tuned. Assume that $\ell^{\mathbf{W},\mathbf{b}}(\Delta\mathbf{W},\widetilde{\mathbf{Z}})$ is $\sigma$-sub-Gaussian2 under $(\Delta\mathbf{W},\widetilde{\mathbf{Z}})\sim P_{\Delta\mathbf{W}\mid\mathbf{W},\mathbf{b}}\times\mu$. Then, for LoRA (adapting $A$ and $B$):
$\bigl{|}\text{gen}\bigl{(}\mu,\mathcal{A}_{\text{BA}}\bigr{)}\bigr{|}\;\leq\;\sqrt{\frac{2q\sigma^{2}\ln 2}{n}\sum_{i\in\mathcal{I}}\Bigl{(}rd_{\text{in}}^{(i)}+rd_{\text{out}}^{(i)}\Bigr{)}},$ (3)
and for truncated SVD adaptation (adapting $B$, $E$, and $A$):
$\bigl{|}\text{gen}\bigl{(}\mu,\mathcal{A}_{\text{BEA}}\bigr{)}\bigr{|}\;\leq\;\sqrt{\frac{2q\sigma^{2}\ln 2}{n}\sum_{i\in\mathcal{I}}\Bigl{(}rd_{\text{in}}^{(i)}+rd_{\text{out}}^{(i)}+r\Bigr{)}},$ (4)
here, $n$ denotes the training set size, and $q$ reflects the number of bits used to encode each dimension in tunable parameters. Due to the low-rank dimension ($r\ll\min(d_{\text{in}},d_{\text{out}})$), truncated SVD adaptation achieves almost the same generalization bound to LoRA.

### v3

@@ Abstract
Pre-trained Language Models (PLMs) have demonstrated their superiority and versatility in modern Natural Language Processing (NLP), effectively adapting to various downstream tasks through further fine-tuning. Federated Parameter-Efficient Fine-Tuning (FedPEFT) has emerged as a promising solution to address privacy and efficiency challenges in distributed training for PLMs on resource-constrained local devices. However, our measurements reveal two key limitations of FedPEFT: heterogeneous data across devices exacerbates performance degradation of low-rank adaptation, and a fixed parameter configuration results in communication inefficiency. To overcome these limitations, we propose FedARA, a novel Adaptive Rank Allocation framework for federated parameter-efficient fine-tuning of language models. Specifically, FedARA employs truncated Singular Value Decomposition (SVD) adaptation to enhance similar feature representation across clients, significantly mitigating the adverse effects of data heterogeneity. Subsequently, it utilizes dynamic rank allocation to progressively identify critical ranks, effectively improving communication efficiency. Lastly, it leverages rank-based module pruning to automatically remove inactive modules, steadily reducing local computational cost and memory usage in each federated learning round. Extensive experiments show that FedARA consistently outperforms baselines by an average of 6.95% to 8.49% across various datasets and models under heterogeneous data while significantly improving communication efficiency by 2.40$\times$. Moreover, experiments on various edge devices demonstrate substantial decreases in total training time and energy consumption by up to 48.90% and 46.95%, respectively.
@@ I Introduction
Pre-trained Language Models (PLMs) [1, 2, 3, 4] have demonstrated extraordinary performance and foundational capabilities in modern Natural Language Processing (NLP). Fine-tuning of PLMs further enhances their performance on downstream tasks, such as sequence classification, summarization, and question answering. Pre-training and fine-tuning are evolving into two distinct yet mutually reinforcing stages in NLP. 1) Pre-training: it involves self-supervised training on large-scale corpora from sources like books, news articles, and Wikipedia, aiming to learn general representations of contextual language and create a generalizable pre-trained model. 2) Fine-tuning: it extends the pre-trained model to specific downstream tasks through supervised training, optimizing its performance to meet domain requirements. Overall, these two stages greatly enhance the model’s generality and flexibility across diverse NLP tasks.
@@
However, the massive scale of PLMs, often with hundreds of millions to billions of parameters [10], incurs substantial computation and communication costs in FL, rendering deployment on resource-constrained devices nearly impossible. To address this issue, a promising solution is to adopt Parameter-Efficient Fine-Tuning (PEFT) methods, which train only a minimal set of parameters using adaptable modules while keeping the pre-trained model frozen [11, 12, 13, 14, 15]. Among these PEFT methods, Low-Rank Adaptation (LoRA) [15] has gained widespread adoption for its effectiveness and efficiency, achieved by introducing two trainable low-rank matrices alongside the pre-trained model for diverse task adaptation. By combining PEFT with FL, distributed data resources on local devices can be fully leveraged for collaborative fine-tuning in a privacy-preserving manner. Consequently, Federated Parameter-Efficient Fine-Tuning (FedPEFT) shows significant advancements and finds wide-ranging applications across various domains [16, 17, 18, 19, 20, 21, 22]. Building on this progress, FedPEFT has increasingly gained recognition as a key method for distributed fine-tuning on local devices [23, 24]. Figure 1 shows the complete FedPEFT training process with LoRA in local clients.
Nevertheless, FedPEFT still faces two major challenges when applied to local devices. 1) FedPEFT exacerbates the impact of non-independent and identically distributed (non-IID) data on model performance. Although performance degradation caused by non-IID data is a well-known challenge in FL, our preliminary measurements show that FedPEFT exacerbates this issue significantly. Specifically, on the 20News [25] dataset using the DistilBERT [2] model, FedPEFT methods exhibit an average accuracy drop of 14.62% under non-IID conditions compared to IID settings—more than double the 6% drop observed with federated full fine-tuning [19]. 2) Fixed rank configurations in FedPEFT incur dual losses in performance and communication cost. Low-rank parameters can be inserted into various positions within the model, such as layers or components like the query, key, and value matrices. However, static allocations often underfit high-sensitivity positions, leading to suboptimal performance—our experiments reveal accuracy gaps of up to 9.76% across different configurations. Meanwhile, redundant parameters assigned to less critical components are repeatedly transmitted during FL rounds without corresponding performance gains, resulting in communication inefficiency.
These dual losses become more severe under non-IID settings, where data heterogeneity intensifies the subspace divergence of federated low-rank adaptation among client-specific models. This divergence leads to greater variability in rank requirements, making fixed rank configurations increasingly misaligned with actual capacity needs. Consequently, both performance degradation and communication overhead are further amplified. Therefore, these two challenges—data heterogeneity and fixed rank allocation—must be jointly addressed to fully unlock the potential of FedPEFT in practical federated learning scenarios.
@@
To address these challenges, we propose Adaptive Rank Allocation for Federated PEFT of language models (FedARA), designed to enhance performance and communication efficiency with minimized resource costs by integrating three novel methods: 1) Truncated SVD adaptation (Section IV-A). We hypothesize that the performance degradation of LoRA arises from its inherent structure and asymmetric initialization. Inspired by Singular Value Decomposition (SVD), we propose a novel truncated SVD adaptation scheme that inserts a diagonal matrix into the low-rank subspace to enable independent feature scaling while preserving symmetric initialization. This mitigates client drift and enhances cross-client feature alignment, significantly alleviating the negative impact of non-IID data. 2) Dynamic rank allocation (Section IV-B). To avoid fixed parameter configurations, we present a dynamic rank allocation method that generates local rank masks through triplets on each client and produces global rank masks via threshold-based arbitration on the server. During federated fine-tuning, less critical ranks are progressively pruned, retaining only significant ones based on global rank masks, effectively improving the communication efficiency of FedPEFT. 3) Rank-based module pruning (Section IV-C). Building on the above two methods, the rank of SVD modules dynamically evolves over FL rounds, we thereby adopt a rank-based module pruning mechanism to automatically remove inactive modules during training. By continuously monitoring rank changes, modules with ranks that fall to zero are safely excluded from the trainable parameters, steadily reducing the computational and storage overhead for client-side training.
@@
Results. We implemented FedARA and evaluated it on three representative edge devices, NVIDIA Jetson AGX Orin, Orin Nano, and Raspberry Pi 5, whose hardware profiles reflect typical resource-constrained settings. We utilize three pre-trained models across five diverse NLP datasets. FedARA outperforms baselines [22, 26, 19, 20] on classification tasks by an average of 6.95% to 8.49% under non-IID data while significantly improving communication efficiency by an average of 2.40$\times$. Remarkably, the total training time on the above devices also decrease by 27.65%, 46.57%, and 48.90%, respectively, and the energy consumption drops by 46.95% on Orin Nano. Ablation experiments reveal that truncated SVD adaptation improves the average accuracy by 7.71%, dynamic rank allocation reduces communication overhead by 70.82% per round while maintaining on-par performance, and rank-based module pruning lowers the average local computation and peak GPU memory usage per round by 10.81% and 31.67%, respectively.
@@
  • We create and integrate three innovative methods into FedARA: 1) truncated SVD adaptation strengthens the structure and adjusts initialization to enhance similar feature representation across clients; 2) dynamic rank allocation generates rank masks locally and arbitrates them globally to reduce the communication of less important parameters; 3) rank-based module pruning monitors dynamic rank changes to remove inactive modules.
@@ II-A Federated Parameter-Efficient Fine-Tuning (FedPEFT)
FedPEFT orchestrates PEFT strategies within the FL framework to efficiently update local models, drastically reducing communication overhead by only transmitting a small number of PEFT modules. These PEFT modules can be replaced by Adapter [11, 12], Prompt[13, 14], or LoRA [15]. Among them, LoRA-based methods are especially popular because they introduce no additional latency during inference. FedPEFT not only preserves privacy but also effectively aggregates diverse knowledge across clients, facilitating the collaborative development of a high-quality global model. By harnessing the advantages of both PEFT and FL, FedPEFT has become an essential technique for deploying fine-tuning on resource-constrained devices [23, 24, 20, 18, 17, 21].
@@ II-B Preliminary Measurements
Observation-1: FedPEFT is more effective than FedFFT on advanced edge devices. FedPEFT comprises two main methods: adapter-based [23] and LoRA-based methods [16, 17, 18, 19, 21]. LoRA-based methods have gradually replaced adapter-based methods to avoid additional inference latency. According to established work [23], FedPEFT achieves comparable accuracy to federated full fine-tuning (FedFFT) while significantly reducing local training time by 1.6$\times$, communication cost by 200$\times$, and memory usage by 1.5$\times$ on previous-generation edge devices like NVIDIA Jetson TX2, Nano, and Raspberry Pi 4B. However, its impact on overall training time on advanced devices remains underexplored. To fill this gap, we evaluate FedLoRA on the 20News [25] dataset with the BERT [3] model using latest edge devices. Following FedPEFT’s typical settings [23], we set the communication bandwidth to 1MB/s. As depicted in Figure 2a, the results indicate that FedFFT’s overall training time is 42$\times$, 40$\times$, and 18$\times$ longer than FedLoRA on AGX Orin, Orin Nano, and RPi 5, respectively. FedPEFT demonstrates its effectiveness compared to FedFFT for cutting-edge embedded devices.
@@
Observation-3: The contribution of PEFT modules in FedPEFT to performance varies their inserted positions, leading to suboptimal performance and inefficient communication in fixed parameter configurations. In FedPEFT, low-rank modules can be placed at various layers and components, such as Query (Q), Key (K), Value (V), Out (O), F1, and F2. Here, O denotes the linear layer in the multi-head attention block, while F1 and F2 represent the two linear layers of the feed-forward network. As shown in Figure 2c, we allocated a fixed rank budget to specific components or layers to evaluate their impact on accuracy. The left sub-figure shows the allocation of the entire rank budget to a single layer, while the right sub-figure illustrates its distribution across a single component in all layers. The findings reveal that deeper layers near the output often have a more significant effect than shallower ones, with components F1 and F2 typically exerting a greater influence than component O. Moreover, the impact of these inserted positions varies across datasets, suggesting that there is no one-size-fits-all configuration. Fixed rank configurations can result in inefficient communication, as they repeatedly transmit less impactful parameters over hundreds of rounds, yielding minimal performance gains under a limited parameter budget.
@@
Insights: These observations underscore key directions for advancing our research. First, given its established advantages, such as improved training efficiency, future research could benefit from moving beyond comparisons with FedFFT and instead focusing on addressing the unique challenges it presents in FL. Second, the performance degradation of different PEFT methods under non-IID remains a critical barrier, necessitating further investigation into effective approaches. Third, fixed parameter configurations lead to communication inefficiency by repeatedly transmitting less impactful parameters over multiple FL rounds, particularly in resource-constrained environments. Finally, optimizing FedPEFT for edge devices requires further refinement, as different hardware presents distinct bottlenecks. To bridge these gaps, enhancing fine-tuning structures and dynamically adjusting parameters could be essential while simultaneously minimizing both communication and computational overhead.
@@ III Overview of FedARA
11:     $\{\hat{P}_{l+1}\},\{M_{l+1}\}\leftarrow\text{CLIENT\_TRAIN}(r,\hat{P}_{g},M_{g})$
@@ IV Design Details of FedARA
In this section, we present the design of FedARA by elaborating on three interrelated methods: truncated SVD adaptation, dynamic rank allocation, and rank-based module pruning.
@@ IV-A Truncated SVD Adaptation
$W=W_{\text{pre}}+\Delta W=W_{\text{pre}}+\frac{\alpha}{r}BA,$ (1)
where $W_{\text{pre}},\Delta W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, $B\in\mathbb{R}^{d_{\text{out}}\times r}$, $A\in\mathbb{R}^{r\times d_{\text{in}}}$, and $\frac{\alpha}{r}$ is a scaling factor with $r\ll\min\{d_{\text{out}},d_{\text{in}}\}$. During training, only the low-rank matrices $A$ and $B$ are updated, while $W_{\text{pre}}$ remains frozen. This decomposition improves parameter efficiency when $r(d_{\text{out}}+d_{\text{in}})<d_{\text{out}}d_{\text{in}}$. Typically, $A$ is initialized with Gaussian noise, and $B$ is initialized to zero, ensuring $\Delta W=0$ at initialization.
@@
To address this issue, we draw inspiration from the structure of truncated Singular Value Decomposition (SVD), which decomposes a matrix into the eigenvalues and eigenvectors of its low-rank subspace. Accordingly, we propose a truncated SVD adaptation method that mirrors this structure, formulated as
$W=W_{\text{pre}}+\Delta W\;=\;W_{\text{pre}}\;+\;\frac{\alpha}{r}BEA,$ (2)
where $E\in\mathbb{R}^{r\times r}$ is a diagonal matrix, and $B\in\mathbb{R}^{d_{\text{out}}\times r}$, $A\in\mathbb{R}^{r\times d_{\text{in}}}$. Both $A$ and $B$ are initialized with random Gaussian values, while $E$ is initialized to zero, ensuring that $\Delta W=0$ at the beginning of training. In contrast to the standard truncated SVD, we do not impose orthogonality constraints on $B$ and $A$, as such constraints reduce flexibility and incur additional computational overhead in each training round.
@@
As illustrated in Figure 4, a diagonal matrix $E$ is inserted between the low-rank subspaces at negligible cost ($r\ll\min(d_{\text{in}},d_{\text{out}})$. The diagonal structure of $E$ restricts updates to independent scaling along each latent dimension, which helps mitigate inter-client variability under non-IID data. At the same time, symmetric initialization ensures balanced gradient flow, offering clients a stable starting point and enhancing global consistency during federated fine-tuning.
We now formally provide theoretical insights into the effectiveness of truncated SVD adaptation under non-IID settings. For each client $k$, let the rank–$r$ update be defined as:
$\Delta W_{k}^{\text{BA}}=\sum_{i=1}^{r}b_{k,i}a_{k,i}^{\top},\quad\Delta W_{k}^{\text{BEA}}=\sum_{i=1}^{r}e_{k,i}b_{k,i}a_{k,i}^{\top},$ (3)
and we quantify client heterogeneity by the drift variance:
$\sigma^{2}=\frac{1}{K}\sum_{k=1}^{K}\!\bigl{\|}\Delta W_{k}-\overline{\Delta W}\bigr{\|}_{F}^{2}\;\approx\;\mathbb{E}\!\left[\!\|\Delta W\|_{F}^{2}\right].$ (4)
Let $T_{i}$ denote the $i$-th rank-1 term. Then, the expression $\|\Delta W\|_{F}^{2}$ can be reformulated as:
$\|\Delta W\|_{F}^{2}=\sum_{i=1}^{r}\sum_{j=1}^{r}\langle T_{i},T_{j}\rangle,$ (5)
where $\langle T_{i},T_{j}\rangle=\operatorname{Tr}(T_{i}^{\top}T_{j})$. Using the identity $\langle uv^{\top},xy^{\top}\rangle=(u^{\top}x)(v^{\top}y)$, the inner product becomes:
$\langle T_{i},T_{j}\rangle=\begin{cases}(b_{i}^{\top}b_{j})(a_{i}^{\top}a_{j}),&\text{(BA)}\\ e_{i}e_{j}\,(b_{i}^{\top}b_{j})(a_{i}^{\top}a_{j}),&\text{(BEA)}\end{cases}\quad,$ (6)
We adopt the standard separable-correlation model:
$\mathbb{E}[b_{i}^{\top}b_{j}]=\begin{cases}\tau_{b},&i=j\\ \rho_{b},&i\neq j\end{cases}\quad,\quad\mathbb{E}[a_{i}^{\top}a_{j}]=\begin{cases}\tau_{a},&i=j\\ \rho_{a},&i\neq j\end{cases}\quad,$ (7)
where $\rho_{a}\rho_{b}\neq 0$ captures non-IID data. For the truncated SVD adaptation, we further assume:
$\mathbb{E}[e_{i}]=0,\quad\mathbb{E}[e_{i}^{2}]=\tau_{e},\quad\bigl{|}\mathbb{E}[e_{i}e_{j}]\bigr{|}\leq\frac{c}{r}\quad(i\neq j).$ (8)
Assumption (8) is justified since each $e_{i}$ is initialized as $e_{i}^{(0)}=0$ and updated by $e_{i}\leftarrow e_{i}-\eta\langle\nabla\ell(x;W),\,b_{i}a_{i}^{\top}\rangle$, where $e_{i}$ only interacts with its own rank-1 component, leading to independent gradient noise across $i\neq j$. Under symmetric mini-batch sampling, the noise is unbiased with $\mathbb{E}[\Delta e_{i}]=0$. Additionally, with Gaussian initialization, the inner product $(b_{i}^{\top}b_{j})(a_{i}^{\top}a_{j})=O(1/\sqrt{d})$, yielding residual correlation of order $O(1/r)$ when $r<\sqrt{d}$.
Substituting equations (6)–(8), we obtain for LoRA:
$\mathbb{E}\!\left[\|\Delta W^{\text{BA}}\|_{F}^{2}\right]=r\tau_{b}\tau_{a}+r(r-1)\rho_{b}\rho_{a}=\Theta(r^{2}),$
while for truncated SVD adaptation, the $r(r-1)$ cross terms are scaled by $\mathbb{E}[e_{i}e_{j}]$ and vanish up to order $O(r)$, yielding:
$\mathbb{E}\!\left[\|\Delta W^{\text{BEA}}\|_{F}^{2}\right]=r\tau_{e}\tau_{b}\tau_{a}+O(r)=\Theta(r).$
Thus, the variance scales as $\sigma^{2}_{\text{BA}}=\Theta(r^{2})$ and $\sigma^{2}_{\text{BEA}}=\Theta(r)$, confirming that truncated SVD adaptation eliminates the quadratic cross-covariance terms and reduces drift growth from $\Theta(r^{2})$ to $\Theta(r)$. This theoretically supports the effectiveness of the newly diagonal matrix $E$ in mitigating non-IID effects.